OpenStudy (anonymous):

What is the y−coordinate of the y−intercept of the line that passes through the points (−4, −4) and (4, 8)?

OpenStudy (anonymous):

it's blue

OpenStudy (anonymous):

its blue?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

what are you talking about?

OpenStudy (anonymous):

Firstly find the slope: \[Slope = \frac{y_2 - y_1}{x_2 - x_1} \implies \frac{12}{8} = \frac{3}{2}\] Let y = mx + c Put any of the point in this to find c: Put y = -4 and x = -4 So: \[-4 = \frac{3}{2}(-4) + c \implies c = -4 + 6 \implies \color{green}{c = 2}\] So this is the y coordinate of y intercept..